# Can a Crow Hatch a Falcon? Lineage Matters in Predicting Large Language Model Performance

### v1

@@ Appendix A Detailed Model Statistics
Figure 1: Distribution of lineage connections among the Qwen 2.5-7B family. The distribution approximates a logarithmic trend.
Figure 2: Complete lineage network visualization among the Qwen 2.5-7B models. The most densely connected node is Qwen-2.5-7B-Instruct.
A.2 Experiment 2: Heterogeneous Models (All Models)
Lineage Analysis of Heterogeneous Models
We also conducted a lineage analysis across the full heterogeneous set, uncovering again a roughly logarithmic distribution in the number of lineage connections (Figure 3). Approximately 60% of these models are either fine-tuned or merged variants, suggesting lineage information is widely applicable for modeling relationships and improving predictions across a broad spectrum of publicly available models.
Table 4: Composition of the heterogeneous set used in Experiment 1.
@@ A.2 Experiment 2: Heterogeneous Models (All Models)
Figure 3: Distribution of lineage connections across the heterogeneous model set. A logarithmic distribution is again apparent, indicating many models have relatively few lineage connections, with only a minority having extensive lineages.
Appendix B Hyperparameter Tuning
Our experiments involve performance data collected from six benchmark tasks provided by the Hugging Face Open LLM Leaderboard, totaling approximately 21,000 instances. For each instance, we have binary outcomes indicating whether a given LLM correctly solved the instance or not. The task is thus formulated as predicting these binary instance-level outcomes for model-instance pairs for which no evaluation is available.
Hyperparameter tuning for all methods was performed using the validation set, optimizing for the instance-level AUC-ROC metric, as this metric effectively captures the method’s capability to distinguish between correct and incorrect model predictions. Below, we detail the hyperparameter search procedures and illustrate the effect of varying the number of observed instances ($t$) on prediction quality.
B.1 Lineage-Regularized MF
@@
For the instance similarity regularization parameter $\lambda_{\mathcal{X}}$, we observe that performance remains relatively stable (AUC-ROC $\approx$ 0.60) for values between $10^{-9}$ and $10^{-6}$, then gradually improves to peak at $\lambda_{\mathcal{X}}\approx 10^{-4}$ with an AUC-ROC of approximately 0.62. Beyond this optimal value, performance decreases significantly, dropping to an AUC-ROC of approximately $0.55$ when $\lambda_{\mathcal{X}}\geq 10^{-2}$.
For the lineage regularization parameter $\lambda_{\mathcal{M}}$, we observe remarkably stable performance across all tested values ranging from $10^{-9}$ to $10^{-3}$, with AUC-ROC consistently around $0.62-0.63$. This stability suggests that model lineage information provides robust guidance for performance prediction across a wide range of regularization strengths, with a slight preference for values at the lower end of the spectrum.
These results indicate that while instance similarity regularization requires careful tuning with an optimal value around $10^{-4}$, the lineage-based regularization is considerably more robust to parameter selection. This finding highlights the inherent stability of lineage information as a guiding constraint for performance prediction.
Once the best hyperparameter values are identified on the development set, we retrain Lineage-Regularized MF on the combined training and development sets and then report the final performance on the test set. The same procedure is applied to all baselines to ensure a fair comparison.
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
@@
Figure 4: Development-set AUC-ROC of Lineage-Regularized MF as a function of the lineage regularization parameter $\lambda_{\mathcal{M}}$ and the instance regularization parameter $\lambda_{\mathcal{X}}$. Each plot fixes one parameter while sweeping the other. Optimal performance is achieved with moderate regularization ($\lambda_{\mathcal{M}}\approx 10^{-4}$).
Once the best hyperparameter values are identified on the development set, we retrain Lineage-Regularized MF on the combined training and development sets and then report the final performance on the test set. The same procedure is applied to all baselines to ensure a fair comparison.
@@
Appendix C Effect of the Number of Observed Instances ($t$)
@@
Figures 5 and 6 present the test-set AUC-ROC across two critical evaluation scenarios:
  • In-distribution prediction: Performance on withheld instance-model pairs where the models appeared in training but certain instances were unobserved (Figure 5(a) and 6(a)).
  • Cold-start prediction: Zero-shot performance on entirely new models unseen during training, representing the challenging cold-start scenario in LLM development (Figure 5(b) and 6(b)).
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
Figure 5: AUC-ROC as a function of observed instances ($t$) per model, evaluated on (a) unobserved instance-model pairs within training models and (b) completely unseen test models for the homogeneous Qwen 2.5-7B family.
(a) Unobserved pairs within training models
(b) Unobserved pairs for unseen test models
Figure 6: AUC-ROC as a function of observed instances ($t$) per model, evaluated on (a) unobserved instance-model pairs within training models and (b) completely unseen test models for the heterogeneous model set derived from the Hugging Face Open LLM Leaderboard.
@@
The performance gap between methods is especially pronounced in the homogeneous setting (Figure 7(b)), where Lineage-Regularized MF achieves with just 50-100 instances per model what baseline methods require 500+ instances to match. This 5-10x reduction in required evaluations translates to substantial computational savings during model development. In the heterogeneous setting (Figure 7(a)), while all methods benefit from the diversity of models, our approach still maintains a consistent advantage across all observation levels.
(a) Heterogeneous Models (All Models)
(b) Homogeous Models (Qwen-2.5-7B)
Figure 7: Correlation between predicted and actual model performance as the number of observed instances per model ($t$) varies. (a) shows results for the heterogeneous set of all models, and (b) for the homogeneous set of models derived from Qwen 2.5-7B. Lineage-Regularized MF maintains consistently high prediction accuracy even with significantly fewer observations, demonstrating robustness compared to baseline methods.

### v2

@@ A.2 Experiment 2: Heterogeneous Models (All Models)
Table 2: Composition of the heterogeneous model set (All models from the Hugging Face Open LLM Leaderboard) used in Experiment 2.
@@
(a) Qwen 2.5-7B family models
(b) All models
Figure 4: Distribution of lineage connections among models. Both distributions approximate a logarithmic trend, indicating many models have relatively few lineage connections, with only a minority having extensive lineages. (a) Homogeneous Qwen 2.5-7B family. (b) All models from the Hugging Face Open LLM Leaderboard.
Figure 5: Lineage network visualization among the Qwen 2.5-7B models. The most densely connected model is Qwen-2.5-7B-Instruct.
Appendix B Hyperparameter Tuning and Ablation Study
Our experiments involve performance data collected from six benchmark tasks provided by the Hugging Face Open LLM Leaderboard, totaling approximately 21,000+ instances. For each instance, we have binary outcomes indicating whether a given LLM correctly solved the instance or not. The task is thus formulated as predicting these binary instance-level outcomes for model-instance pairs for which no evaluation is available.
B.1 Experimental Protocol
All hyperparameter optimization was performed on the validation set using two complementary metrics:
  • Primary metric: Instance-level AUC-ROC, which captures the model’s ability to distinguish between correct and incorrect predictions
  • Secondary metric: Pearson correlation coefficient, which measures how well the method preserves relative performance superiority relationship among models
For all methods, we performed grid searches over relevant regularization parameters and selected the configuration that maximized AUC-ROC on the development set. Additionally, we conducted comprehensive ablation studies using Pearson correlation to better understand the contribution of each component.
B.2 Lineage-Regularized MF
@@ B.2.1 AUC-ROC Analysis
For the model lineage regularization parameter $\lambda_{\mathcal{M}}$, we observe a clear improvement as the regularization strength increases, with AUC-ROC rising from approximately 0.795 at $\lambda_{\mathcal{M}}=10^{-9}$ to a peak of approximately 0.895 at $\lambda_{\mathcal{M}}=10^{-4}$, demonstrating the critical importance of lineage information for performance prediction.
@@
For the instance similarity regularization parameter $\lambda_{\mathcal{X}}$, we observe that performance remains highly stable (AUC-ROC $\approx$ 0.895) for values at $10^{-5}$, and then decreases to approximately 0.86 when $\lambda_{\mathcal{X}}\geq 10^{-3}$, indicating over-regularization.
These results indicate that model lineage regularization ($\lambda_{\mathcal{M}}$) has a more pronounced impact on performance, with optimal values around $10^{-4}$, while instance similarity regularization ($\lambda_{\mathcal{X}}$) shows broader stability across multiple orders of magnitude.
(a) Varying $\lambda_{\mathcal{M}}$ (model lineage) with fixed $\lambda_{\mathcal{X}}=10^{-5}$
(b) Varying $\lambda_{\mathcal{X}}$ (instance similarity) with fixed $\lambda_{\mathcal{M}}=10^{-4}$
@@
Figure 6: Development-set AUC-ROC of Lineage-Regularized MF as a function of the lineage regularization parameter $\lambda_{\mathcal{M}}$ and the instance regularization parameter $\lambda_{\mathcal{X}}$. Each plot fixes one parameter while sweeping the other. Optimal performance is achieved with moderate regularization ($\lambda_{\mathcal{M}}\approx 10^{-4}$, $\lambda_{\mathcal{X}}\approx 10^{-5}$, ).
B.2.2 Comprehensive Ablation Study
To more thoroughly understand the interaction between regularization components and their individual contributions, we conducted an extensive ablation study measuring Pearson correlation coefficients across the full hyperparameter space (Figure 7).
Critical observations from the ablation study:
  1. 1.
    Baseline failure in cold-start: When both $\lambda_{\mathcal{M}}=0$ and $\lambda_{\mathcal{X}}=0$ (equivalent to standard Matrix Factorization/NCF), the method completely fails in cold-start scenarios with correlation $\approx 0$, confirming that vanilla collaborative filtering cannot handle unseen models.
  2. 2.
    Model lineage as the dominant factor: Setting $\lambda_{\mathcal{M}}>0$ while keeping $\lambda_{\mathcal{X}}=0$ achieves Pearson correlation coefficients of 0.20.20.50.5, demonstrating that model lineage alone provides substantial predictive power even without instance similarity information. This shows that most of the correlation improvement stems from the model lineage regularization term.
  3. 3.
    Synergistic effects: The optimal configuration ($\lambda_{\mathcal{M}}\approx 10^{-4}$, $\lambda_{\mathcal{X}}\approx 10^{-5}$) achieves correlation >0.5>0.5, with the heatmap revealing a clear “sweet spot” (highlighted by the red box) where both regularization terms work synergistically.
  4. 4.
    Robustness of model lineage: The correlation remains stable across a wide range of $\lambda_{\mathcal{M}}$ values ($10^{-6}$ to $10^{11}$), indicating that lineage information is inherently robust and does not require precise tuning.
Figure 7: Heatmap of Pearson correlation coefficients for all combinations of $\lambda_{\mathcal{M}}$ (model lineage) and $\lambda_{\mathcal{X}}$ (task/instance similarity) regularization parameters. The red box highlights the optimal region. Note that $\lambda_{\mathcal{M}}=0,\lambda_{\mathcal{X}}=0$ (bottom-left) corresponds to standard MF and fails completely in cold-start prediction (correlation $\approx 0$). The visualization clearly shows that model lineage regularization ($\lambda_{\mathcal{M}}>0$) is the primary driver of performance improvement.
These ablation results validate our core hypothesis that explicit lineage relationships provide crucial signals for performance prediction, with the model lineage regularization term contributing the majority of the improvement. Once the best hyperparameter values are identified on the development set, we retrain Lineage-Regularized MF on the combined training and development sets and report the final performance on the test set.
B.3 Ablation: Instance-graph construction
Embedding backbones.
We evaluated several widely used embedding models for constructing the instance-similarity graph—Snowflake/snowflake-arctic-embed-l-v2.0, infloat/e5-mistral-7b-instruct, and all-mpnet-base-v2. For each model, we embedded the prompt text, computed cosine similarities, and formed a $k$-NN graph. Across benchmarks, we observed no material differences in downstream metrics (AUC-ROC and Pearson correlation), and the relative ordering of methods remained stable. Given this robustness and for consistency, we use Snowflake/snowflake-arctic-embed-l-v2.0 in the main results.
Neighborhood size $k$.
We swept $k\in\{2,5,10,20,50,100\}$ on the development set, selecting $k$ by AUC-ROC. Performance was largely insensitive to $k$ across tasks; we therefore fix k=20k{=}20 in all reported experiments as a stable sparsity–connectivity trade-off.
Appendix C Robustness to Lineage Noise
To quantitatively assess the impact of noise/incompleteness in lineage data, we conducted additional experiments by randomly modifying the lineage information (either adding or removing lineage links) in the heterogeneous setting. Notably, we observed distinct responses between the two methods:
  • Lineage-Regularized MF: Robust to incomplete (missing) lineage data (removal of 40% of links led to only a 10% decrease in correlation), but highly sensitive to incorrect lineage additions (adding 40% false links caused a 50% decrease in correlation). For LRMF, accuracy is better preserved by omitting uncertain lineage information rather than risking incorrect additions.
  • Model Lineage Average: Demonstrated robustness against both random additions and removals of lineage data. This resilience likely stems from random lineage connections averaging out model-specific variations and pulling results toward a global mean.
Figure 8: Effect of noise in lineage data on model prediction accuracy in the heterogeneous setting, measured by Pearson correlation. Noise is introduced either by randomly removing lineage links (negative values) or randomly adding spurious links (positive values). LRMF shows robustness to missing links but suffers significant degradation when incorrect lineages are added, while MLA remains comparatively stable under both scenarios.
@@ Appendix C Robustness to Lineage Noise
Appendix D Impact of the Number of Observed Instances per Model on Prediction Performance
@@
Figures 9 and 10 summarize the test-set AUC-ROC across two critical evaluation scenarios for both the homogeneous (Qwen 2.5-7B family) and heterogeneous (all models) settings:
  • In-distribution prediction: Estimating performance on unobserved instance-model pairs, where the models themselves are included in the training set but certain instances remain unobserved (shown in Figures 9(a) and 10(a) for the homogeneous and heterogeneous settings, respectively). While this differs from the main cold-start scenario, it requires imputing missing data due to limited observations per model.
  • Cold-start prediction: Zero-shot performance on entirely new models unseen during training, representing the main cold-start scenario of this work (shown in Figures 9(b) and 10(b) for the homogeneous and heterogeneous settings, respectively).
(a) In-distribution prediction: AUC–ROC for withheld instance–model pairs within training models
(b) Cold-start prediction: AUC–ROC for instance–model pairs of unseen test models
Figure 9: AUC–ROC as a function of the number of observed instances per model ($t$) in the homogeneous Qwen 2.5-7B family: (a) in-distribution prediction and (b) cold-start prediction.
(a) In-distribution prediction: AUC–ROC for withheld instance–model pairs within training models
(b) Cold-start prediction: AUC–ROC for instance–model pairs of unseen test models
Figure 10: AUC–ROC as a function of the number of observed instances per model ($t$) in the heterogeneous Hugging Face model set: (a) in-distribution prediction and (b) cold-start prediction.
@@
The performance gap between methods is especially pronounced in the homogeneous setting (Figure 11(a)), where Lineage-Regularized MF achieves with just 50-100 instances per model what baseline methods require 500+ instances to match. This 5-10x reduction in required evaluations translates to substantial computational savings during model development. In the heterogeneous setting (Figure 11(b)), while all methods benefit from the diversity of models, our approach still maintains a consistent advantage across all observation levels.
(a) Homogeneous setting (Qwen 2.5-7B family): Pearson correlation for cold-start prediction
(b) Heterogeneous setting (all models): Pearson correlation for cold-start prediction
Figure 11: Pearson correlation coefficient for cold-start prediction as a function of the number of observed instances per model ($t$): (a) homogeneous Qwen 2.5-7B family and (b) heterogeneous Hugging Face model set.
@@
Appendix E Routing Details
Below we provide a detailed analysis of the routing assignments produced by each method. For both the heterogeneous model set (all models) and the homogeneous Qwen-2.5-7B family, we report:
  1. 1.
    The top-20 models most frequently selected across all test instances (Figures 1822, Figures 1216).
  2. 2.
    The top-20 models most frequently selected per benchmark (Figures 1923, Figures 1317).
In both the homogeneous (Qwen-2.5-7B family) and heterogeneous (all models) settings, we observe the following common routing behaviors:
LRMF-Based Routing.
LRMF consistently leverages lineage information to select a diverse set of high-performing variants. When aggregated across all tasks, instruction-tuned and merge-derived (“soup”) models dominate the top selections. On a per-benchmark basis, the method dynamically adapts—for example, favoring instruction-tuned models on IFEval and large base models on reasoning benchmarks such as BBH and MMLU-Pro.
Model Lineage Averaging–Based Routing.
The kNN-style lineage averaging approach concentrates almost exclusively on immediate ancestor checkpoints. It shows minimal variability across benchmarks, typically defaulting to whichever parent model achieved the highest raw score.
NCF with Factors–Based Routing.
The factor-based MLP exhibits an overwhelming bias toward a single model across both settings. This behavior indicates overfitting to dominant metadata signals and a lack of task-specific discrimination.
E.1 Homogeneous Qwen-2.5-7B Family setting
E.1.1 LRMF-Based Routing
Figure 12: Top 20 most frequently routed-to models by LRMF-based routing in the homogeneous experiment.
Figure 13: Top 10 most frequently routed-to models by LRMF-based routing in the homogeneous experiment. This demonstrates that this routing approach assigns different models as optimal destinations depending on the benchmark.
E.1.2 Model Lineage Averaging–Based Routing
Figure 14: Top 20 most frequently routed-to models by MLA-based routing in the homogeneous experiment.
Figure 15: Top 10 most frequently routed-to models by MLA-based routing in the homogeneous experiment. This demonstrates that this routing approach assigns different models as optimal destinations depending on the benchmark.
E.1.3 NCF with Factors–Based Routing
Figure 16: Top 20 most frequently routed-to models by NCF with factors-based routing in the homogeneous experiment. This result indicates that the routing strategy assigns all tasks exclusively to two models.
Figure 17: Top 10 most frequently routed-to models by NCF with factors-based routing in the homogeneous experiment. This result indicates that the routing strategy assigns all tasks exclusively to two models.
E.2 Heterogeneous Models setting
E.2.1 LRMF-Based Routing
Figure 18: Top 20 most frequently routed-to models by LRMF-based routing in the heterogeneous experiment.
Figure 19: Top 10 most frequently routed-to models by LRMF-based routing in the heterogeneous experiment. This demonstrates that this routing approach assigns different models as optimal destinations depending on the benchmark.
E.2.2 Model Lineage Averaging–Based Routing
Figure 20: Top 20 most frequently routed-to models by MLA-based routing in the heterogeneous experiment.
Figure 21: Top 10 most frequently routed-to models by MLA-based routing in the heterogeneous experiment. This demonstrates that this routing approach assigns different models as optimal destinations depending on the benchmark.
E.2.3 NCF with Factors–Based Routing
Figure 22: Top 20 most frequently routed-to models by NCF with factors-based routing in the heterogeneous experiment. This result indicates that the routing strategy assigns all tasks exclusively to one model.
Figure 23: Top 10 most frequently routed-to models by NCF with factors-based routing in the heterogeneous experiment. This result indicates that the routing strategy assigns all tasks exclusively to one model.